Thursday, April 6, 2017 Alert downgrade to red, N stable.Black alert at 504, black grid ratio alert, N danger The above is a coded messages for the Resistance Movement as communicated by Cobra on Portal 2012. These messages are not meant for the general population, and are not meant to be understood by us. This article (Message for the Resistance Movement: April 5, 2017)…
Sunday, April 2, 2017
Security breach deflected, evaluate SIMORPH / VTXMORPH
via — The Portal
Thursday, February 2, 2017
Pandora in progress, EELA in progress. 10k96M detected. 96Alters detected, removal in progress.HVBN substable to stable. M minimum requirements met.
Cobra in progress, Pandora in progress, multiple EXMOSS / EELA / ZEOLITE sequences in progress, BIOCHIP layer removal in progress, U96 / TOPAZ activated, ATLANTIS12 clear. Several systems / Isidic security breaches, all deflected, HVBN unstable towards stable. Two PB removal attempts with results pending. M minimum requirements met.